OpenStudy (anonymous):

Find the rate of change for the graph. How did you find the rate of change? I will link the graph.

OpenStudy (nincompoop):

link your entire semester of homework also

OpenStudy (charlotte123):

LOL NIN

OpenStudy (charlotte123):

Laughing out loud sir

OpenStudy (tester97):

oh ninny :P

OpenStudy (bibby):

Look 'ma no hands while I'm violating the Code of conduct and posting irrelevant stuff on some question

OpenStudy (happinessbreaksbones):

∖(∖Huge ghey∖)

OpenStudy (happinessbreaksbones):

T-T didnt work

OpenStudy (ashleyisakitty):

your question jsut god f'd in the b

OpenStudy (nincompoop):

that's gotta be difficult without tick marks

OpenStudy (shamil98):

Rate of change is the slope \[\Huge \frac{ \Delta y }{ \Delta x } = \frac{ vertical~change }{ horizontal~change } = \frac{ rise }{ run }\]

OpenStudy (tester97):

wow shammy actually answered da q cudos to u shammy

OpenStudy (shamil98):

\[\Huge \frak Rate~of~change = ~slope\] dis font tho

OpenStudy (anonymous):

Woah this is complicated

OpenStudy (anonymous):

Can anyone look at the graph and help me? PLeaseeeee :'~ (

OpenStudy (anonymous):

\[Slope=m= \frac{ \Delta y }{ \Delta x } = \frac{ y _{2} - y _{1}}{ x _{2} - x _{1} }= \frac{ rise }{ run }\]

OpenStudy (anonymous):

@Lethal I already know the equation. I just don't know what numbers to put in it from the graph.

OpenStudy (anonymous):

That's going to be difficult without the tick marks

OpenStudy (anonymous):

But to get you started you know it goes through the origin so one ordered pair you could use is, (0,0)

OpenStudy (anonymous):

What other could I use? @Lethal

OpenStudy (anonymous):

I couldn't tell ya even if I had perfect vision

OpenStudy (anonymous):

@Ryan2416 @ranga @Rina.r @RadEn @bookworm00981 @brando1996 @bibby @francoisderozay @Preetha @LadyM @Kittta @shamil98 @dan815 @Darrius @doulikepiecauseidont @Directrix Can anyone help?

OpenStudy (anonymous):

Dude. No one can really help. There is no tick marks, otherwise we'd be done with this question. Now, is there anything else you were provided other than the picture of this ungrateful graph?

OpenStudy (anonymous):

If you can't do then I will find someone else who can so hop off. And that picture shows up just fine on my computer.

OpenStudy (anonymous):

Try (4,25), they seem to match if you put a pen to the graph at 25 and see where it touches the graph

OpenStudy (anonymous):

Ok!

OpenStudy (anonymous):

What'd you get?

OpenStudy (anonymous):

It was (3,25) , but thanks anyways. @doulikepiecauseidont